OpenStudy (anonymous):

I REALLY DON'T UNDERSTAND THIS ASSIGNMENT You will start with four different numbers between 1 to 10. Combine these integers, using each number once, to make an expression equal to 1. Then repeat, using each number once, to make an expression equal to 2, and so on. You'll use the four numbers to make expressions equal to the numbers 1-10, and each expression you make is worth 1 point. For example, say you are given 1, 2, 3, and 4, you could create the following expressions: 1 = (2 – 1) • (4 - 3) 2 = (2 • 1) • (4 - 3) 3 =4 - 3 /(2 + 1) You can combine your numbers to make two- or three-digit numbers. You can use parentheses, and you can even have nested parentheses. You can use exponents. If you use division, the integers should be evenly divisible. For example, 4 ÷ 2 is okay, but 3 ÷ 2 is not okay When you turn in your solution to this problem, make sure that you write down the four numbers that you started with. You may choose any of the following sets of four numbers for this problem: 1, 4, 6, and 8 3, 3, 4, and 5 4, 6, 7, and 8 2, 4, 6, and 8

OpenStudy (sloppycanada):

From what I'm thinking is you're trying to get 1, 4, 6 or 8 or whatever numbers you chose to equal the number 1.

OpenStudy (sloppycanada):

You can only use each number once and you have to use every number.

OpenStudy (sloppycanada):

Your equations must result in full numbers. So if again, you chose 1, 4, 6, and 8, you couldn't divide 6/4. But you could do 8/4.

OpenStudy (anonymous):

I know i have to make expressions for each number b/w 1-10 , but what are the number are the bottom for?

OpenStudy (anonymous):

at*

ganeshie8 (ganeshie8):

This is a fun problem

ganeshie8 (ganeshie8):

Start by picking your favorite four digits

OpenStudy (anonymous):

4,3,1 and 8

ganeshie8 (ganeshie8):

Okay lets combine them and produce a 1

OpenStudy (anonymous):

8-4 + 1-3 =2

ganeshie8 (ganeshie8):

Awesome!

ganeshie8 (ganeshie8):

1*8 - 4 - 3 = 1

OpenStudy (anonymous):

so do each number 1-10 like that?

ganeshie8 (ganeshie8):

yes it may get tricky

ganeshie8 (ganeshie8):

lets see who finds the combination for 3 first

ganeshie8 (ganeshie8):

looks very tricky hmm

ganeshie8 (ganeshie8):

Gotcha!

ganeshie8 (ganeshie8):

8/4 + 1^3 = 3

OpenStudy (anonymous):

1+ 2 + 4 - 3= 4

ganeshie8 (ganeshie8):

but 2 is not ur favorite digit

ganeshie8 (ganeshie8):

we can't use 2 right

OpenStudy (anonymous):

i thought we just make a expressions to equal it , we have to use the same numbers to equal each number 1to 10?

ganeshie8 (ganeshie8):

yes we're allowed to use only the four digits : 4,3,1 and 8

ganeshie8 (ganeshie8):

we need to combine those four digits using the known operations and produce the digits

ganeshie8 (ganeshie8):

are the rules of game clear ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

maybe read the problem again

OpenStudy (anonymous):

my teacher wanted me to redo the assignment , i think i got it now . thanks

ganeshie8 (ganeshie8):

np :) please do share your answer once you have it..